x f
5 5
15 7
25 11
35 16
45 13
55 7
65 4
75 2

mean deviation,variance, standard deviation, coefficent of variance
1
Expert's answer
2011-06-07T05:05:30-0400
m = mean = sum(1:8; f/65 * x) = 36.077
MD = sum(1;8; f /65 * |x-m] ) = 13.6
s = Std deviation = sqrt( sum(1;8; f/65 * |x-m]2 )) = 17.198
Coefficent of variance = s/ |m| = 0.4767